OpenStudy (anonymous):

will fan and medal Which of the following options is an equivalent function to f(x) = 4(3)^2x? both the two and the x are up f(x) = 36x f(x) = 4(9)x f(x) = 144x f(x) = 4(6x)

OpenStudy (anonymous):

f(x) = 36^x f(x) = 4(9)^x f(x) = 144^x f(x) = 4(6x)

OpenStudy (jadedry):

You can't (cleanly) multiply 4 by 3^2x 4 * 3^2x = 4 * 9^x Therefore the answer is 4 *( 9 )^x

OpenStudy (anonymous):

than you